OpenStudy (anonymous):

4x+y=6 y=11x tell whether you would use substitution or elimination method

jimthompson5910 (jim_thompson5910):

y is completely isolated in equation 2, so I'd use the substitution method

jimthompson5910 (jim_thompson5910):

4x+y=6 4x+11x=6 ... substitute in 11x for y (since y = 11x) now solve for x and tell me what you get

OpenStudy (anonymous):

x=2/5? not sure horrible at math

jimthompson5910 (jim_thompson5910):

good, x = 2/5 is correct

OpenStudy (anonymous):

YES!

jimthompson5910 (jim_thompson5910):

now that you know that x = 2/5, you can replace x with 2/5 in either equation that has x,y the second one is easier

jimthompson5910 (jim_thompson5910):

y = 11x y = 11(2/5) y = (11/1)*(2/5) y = (11*2)/(1*5) y = 22/5

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

im new to open study so is there a way i can follow you or something so that you could tutor me again if thats ok?

jimthompson5910 (jim_thompson5910):

you can fan me if you want also, by posting questions to the left, you allow anyone to help and tutor you

OpenStudy (anonymous):

k thx

jimthompson5910 (jim_thompson5910):

sure thing

jimthompson5910 (jim_thompson5910):

also, you can private message any user some users require that you have to be a fan first, but I don't have that restriction so you can private message me if you're a fan or not

jimthompson5910 (jim_thompson5910):

and finally, you can tag users by sticking an @ symbol before their user name to get their attention to come to the post example: @PR0JECTxIC0N

OpenStudy (anonymous):

oh ok thanks again and i fanned you anyway

jimthompson5910 (jim_thompson5910):

alright great, thanks